OpenStudy (anonymous):

The mean diameter of the Sun is 109 times the mean diameter of Earth. Then the volume of the Sun is how many times the volume of EArth?

OpenStudy (anonymous):

(109/2)^3

OpenStudy (anonymous):

the ratio of the radii, cubed

OpenStudy (anonymous):

Close. it's actually 109^3 times the volume of earth, or1295029

OpenStudy (anonymous):

ohhh so all you would have to do for that one is do 109^3 since it's volume?

OpenStudy (anonymous):

Vsphere= 4/3 pi r ^3

OpenStudy (anonymous):

so then would it be 12441.75?

OpenStudy (anonymous):

sun volume = (4/3) pi (109*earth radius)^3 earth volume = (4/3) pi (earth radius)^3 how many times the volume of earth= sun volume/earth volume substitute: how many times the volume of earth = (4/3) pi (109*earth radius)^3/(4/3) pi (earth radius)^3 cancel like terms: how many times the volume of earth= (109*earth radius)^3/ (earth radius)^3 how many times the volume of earth how many times the volume of earth= 109^3*earth radius^3/ earth radius^3 how many times the volume of earth = 109^3

OpenStudy (anonymous):

my mistake, if the diameter is 109 times the diameter of the earth, the radius is also 109 times the radius of the earth